### v2

# Elastic Softening in Synthetic Diamonds

Abstract
This study reveals a novel phenomenon demonstrating the softening of synthetic diamonds when cooled to very low temperatures below 1 K. Herein, we argue that this elastic softening can be attributed to the effect of electric-quadrupole degrees of freedom of the dangling bonds in the neutral single-atom vacancies of carbon. We present the results of ultrasonic investigations of single-crystalline synthetic diamonds, namely type-IIa (colorless) and Ib (yellow) diamonds grown by high-pressure-high-temperature synthesis as well as type-IIa diamond grown by chemical vapor deposition. We observe a magnetic-field-insensitive softening of the elastic constant in all samples at low temperatures below 1 K. Our results strongly suggest a ppb level concentration of neutral single-atomic vacancies in all investigated diamonds. Our findings open new avenues for the quantitative determination of single neutral vacancies in non-irradiated diamonds, an important information needed for their potential application for quantum technology and next-generation semiconductor devices.
Since ancient times, impurities and defects in diamonds have been considered as crucial criteria for appraising jewelry. Recently, the development of synthesis methods for artificial diamonds is of interest not only for semiconductor industry [1], but also for the study of the quantum properties of defects and possibilities for potential applications [2, 3, 4, 5]. Particularly, quantum states of defects arising from atomic vacancies in diamonds are considered promising for application in solid-state quantum sensors in the broad range from life to materials science [6, 7, 8, 9]. Such states also pave the way from fundamental physics to quantum technology, including quantum computation and quantum communication [10, 11, 12, 13, 14, 15]. Quantum-technology studies based on diamonds have primarily utilized commercially available synthetic diamonds prepared through high-temperature-high-pressure (HPHT) or chemical vapor deposition (CVD) methods. Although numerous researchers have investigated how defect centers originating from atomic vacancies with adjacent nitrogen atoms, called NV centers, can be grown and engineered [16, 17, 18, 19], studies comparing and evaluating the defect-center states in diamonds are scarce. In particular, whether the synthetic diamonds used as substrates for quantum devices contain atomic vacancies and to which amount has not yet been explored.
During the synthesis process, the generation of single-atom vacancies on the ppm to ppb level is thermodynamically inevitable [20] and should be well controlled in quantum technologies. Previous studies have primarily reported ppm-level concentrations of neutral single vacancies V on irradiated diamond samples [21, 22, 23]. However, the details of the existence and density of such vacancies (shown in Fig. 1) in non-irradiated synthetic diamonds have not yet been elaborated. Thus, reliable determination of V concentrations at the ppb level and investigation of vacancy properties are crucial for the rational development of novel quantum technologies. Here, we report the successful determination of the neutral single vacancy concentration in the order of ppb by quantitative analysis of the elastic softening based on the quadrupolar susceptibility [20]. Herein, the elastic “softening” refers to a decrease in the elastic stiffness constant. We further demonstrate that ultrasonic measurements at very low temperatures are a robust tool for detecting and studying neutral single vacancies in single-crystalline diamond.
Diamond has a cubic crystal structure (O, , No.227) with lattice constant Å [24], as schematically shown in Fig. 1 (top). The orbitals of carbon with the outermost electron configuration covalently bond to each other and thereby form the diamond structure, resulting in one of the hardest materials on earth. Diamonds exhibit different colors depending on the amount and type of impurities in the crystal. A yellow color is indicative for dominant nitrogen impurities (type Ib), whereas blue-colored diamond indicates the presence of boron as primary impurity (type IIb). There exist two types of colorless diamonds: type Ia, which contains nitrogen atoms in dimer form, and type IIa, which contains only a few impurities. HPHT flux and CVD methods are utilized for synthesizing type-IIa crystals. Recently, the technologies for synthesizing large high-quality single-crystalline diamonds have been significantly improved.
In an earlier study, Goto and coworkers successfully detected V vacancies [25] in single-crystalline Si through ultrasound. Indeed, Goto and coworkers observed a low-temperature softening for some acoustic modes in Si at cryogenic temperatures [20, 26, 27, 28, 29, 30, 31, 32]. Despite significant advancements of crystal-growth techniques and of the electronic properties of large-diameter Si wafers in semiconductor industry, the evaluation of quantum properties at cryogenic temperatures and the neutral single-atomic vacancy concentration leading to these properties remains to be understood properly. This is also true for diamonds. The determination of residual atomic vacancies at the few ppb level, which were previously unattainable (and overlooked) optically, is critical to improve the coherence time of quantum-information devices when downsizing quantum circuits in the future. Therefore, it is crucial to investigate the properties and establish an evaluation procedure for the low vacancy concentration in diamonds for quantum applications.
We used three commercially available synthetic diamond species grown by HPHT synthesis (Sample 1 and 2) and CVD (Sample 3) for our ultrasonic measurements presented here (see SM [20] for experimental details). The temperature dependence of the elastic constant at zero magnetic field is shown in Figs. 2(a) and 2(c) for single-crystalline HPHT type-IIa diamond (Sample 1). gradually hardens down to K [as shown in the inset of Fig. 2(c)], which is a general behavior of solids owing to the anharmonicity of acoustic phonons. This change is consistent with previously reported temperature dependences of the sound velocity of transverse ultrasonic waves down to K [33, 34, 35]. Notably, to the best of our knowledge, no previous work has yet studied the elastic constants below 10 K. The low-temperature region of exhibits 125 ppm softening from 1 K down to 20 mK. In analogy to the interpretation for a similar softening observed in Si, the molecular orbitals of the neutral single vacancies remain at a triple-degenerated quantum state at least down to these temperatures. These results evidence that the local cubic symmetry around a single neutral vacancy V in diamond also remains preserved. In the Sec. 10 of the SM [20], we elaborate the reasons why the effect of NV (as shown in Fig. 1), NV (with electrons removed from NV), large multi-atom vacancy V clusters [36, 37], P centers (neutral single-substitutional nitrogen) [38], and nuclear spins have no effect on the elastic softening.
The solid red line in Fig. 2(a) is the result of calculations based on the equation (for details, see SM [20]). The fit yields mK, mK, and mK. The dashed line represents the phenomenological fit for the background contribution [39], with J m, , and = 2.941652. The negative value of indicates an antiferro-type inter-vacancy interaction for the charge state of the vacancy V. The negative value of suggests quadrupolar fluctuations associated with the triple-degenerate quantum state even at the lowest temperature of 20 mK. This result further proves that the cubic site symmetry at the vacancy site is preserved in the investigated diamond material, i.e., the local distortion owing to the static Jahn-Teller effect is irrelevant for this diamond with low vacancy concentration.
From optical measurements [25, 40, 41] on irradiated diamonds [42], an (vibronic) ground state has been proposed for V [43, 44]. This ground state cannot explain the present elastic softening in . Another experiment on non-irradiated diamond using circular-polarized light recently revealed an symmetry for the ground state of the neutral single atomic vacancy [45]. Therefore, we argue that the ground state of the neutral single vacancy (V) in the non-irradiated diamond has symmetry. A consistent picture of the quantum ground state has not been unambiguously established yet, see Sec. 9 in the SM [20].
Figure 1 illustrates schematically the crystal structure of diamond (top) and the two types of dangling bonds created by atomic vacancies (middle). There are eight atoms in the unit cell, and the atomic positions of the sites are [0,0,0], [,,0], [0,,], and [,0,]. The sites are translationally shifted by [, , ] from those. Both exhibit the same site-symmetry , where the spatial inversion symmetry is broken. Adjacent atoms are located at the vertex and center of the tetrahedron and belong to the crystallographically different and sites, respectively. As the spin-orbit interaction is small in diamond, we do not consider the effect of odd-parity multipoles but only that of even-parity ones [46]. When the molecular orbitals are placed in an electrostatic potential of site symmetry, it results in an energy splitting between a non-degenerate bonding orbital, called singlet and 3-fold degenerate antibonding orbitals, called triplet , as shown schematically at the bottom of Fig. 1 (for corresponding wave functions, see Sec. 5 in the SM [20]).
Moreover, the charge state of the atomic vacancy is determined by the site symmetry of the vacancy and number of electrons in the orbital. To describe the quantum ground state of a single neutral vacancy V, we consider a quantum state in which two electrons occupy the singlet and the other two enter the triplet state. The product space in the triplet state of the group is irreducibly decomposed as , including 9 degrees of freedom. At , the electric quadrupole degrees of freedom belonging only to the and symmetry are active. The quadrupole degrees of freedom that contribute to the softening of are , , and , which exhibit symmetry. The specific matrix elements of the electric quadrupoles can be represented using a Gell-Mann matrix and generated by applying the rotation group SU(3) to the three-dimensional representation [27].
Further, in an NV center the vacancy site exhibits lower symmetry, because of a local symmetry reduction owing to the substitution of one C site by N. Therefore, the orbital splits into two singlets and one doublet [47, 48]. When the nitrogen donor in the substituted position provides the carrier for the valence -1, the NV center is negatively charged (NV). According to Hund’s first rule, the ground state adopts the highest value of total spin , which is consistent with the Pauli exclusion principle. The six electrons occupy the singlets and the and orbits of the doublet to satisfy the condition that all wavefunctions are antisymmetric. Therefore, the ground state of NV results in an state as shown Fig. 1. However, no electric quadrupole degrees of freedom exist in the product space of with . Therefore, the orbitals of NV do not contribute to the softening of .
We estimate the vacancy concentration using the relation for the acoustic mode , which was used to analyze the vacancy concentration in Si as well [27]. Here, denotes the atomic vacancy concentration, the Jahn-Teller energy, and the deformation binding energy. The softening of Czochralski (CZ)-grown Si starts below approximately 10 K, whereas that of diamond starts below about 1 K. In line with that, obtained from the present analysis for diamond is approximately one-tenth of that of Si. Our experiments yield the parameters and , but not , which is unknown for diamond. As a first try, we assume the same value as for Si, i.e., K [29], despite the difference in the dielectric constant, melting temperature, and lattice constant [27]. The estimated neutral single-atomic vacancy concentration for the present sample of type-IIa diamond is 0.2 ppb, exhibiting the same order of magnitude as that for CZ-grown Si, namely 0.11 ppb. Notably, the large difference in the elastic constants, which is seven times larger for diamond than for Si, was also considered in these calculations. and have an error of approximately 5%. Therefore, the parameter is dominant in determining the order of .
If we use the formation energy 8.373 eV of the neutral vacancy, recently estimated by ab-initio calculation [44], we obtain a vacancy concentrations of 0.5 ppb, which is close to our above estimate of 0.02 ppb. Of course, precise values are not meaningful, but we can estimate the order of magnitude depending on how the binding constant is considered. However, there is no doubt that the vacancy concentration is definitely in the sub-ppb level. It would be highly desirable to refine the unknown parameter by theory and ultrasonic experiments using diamonds with known higher vacancy concentration, which can be controlled by irradiation.
According to the theory proposed in a previous study for Si [20, 27], the deformation binding energy can be expressed by a relationship using the quadrupole-strain coupling constant . This coupling constant is proportional to the effective charge and the square of the effective radius of the vacancy orbital (see SM) [20, 49]. In case of Si, a triple-degenerate vacancy state within the band gap can be significantly extended over a large distance of 20 Å around the vacancy site [50]. In conjunction with such theoretical predictions, ultrasonic experiments on boron-doped Si have estimated the value of the strain coupling constant , corresponding to symmetry, to be approximately K [29]. Even if the atomic vacancies are diluted, the spatially extended molecular orbitals of Si atomic vacancies contribute to the reciprocal temperature dependence of the quadrupole susceptibility, which results in the elastic softening at low temperatures. Importantly, the substantial value of allows for the detection of atomic vacancies at the sub-ppb level by use of ultrasound techniques.
Further, we investigated the elastic response for other diamonds grown by different methods. Figure 3 shows the temperature dependence of for three different types of diamonds. Sample 1 is the type-IIa HPHT diamond with data already shown in Fig. 2. Sample 2 is also a type-IIa diamond grown by CVD. Sample 3 is a HPHT yellow diamond of type Ib that presumably contains numerous nitrogen impurities. Sample 1 and 2 show a nearly identical low-temperature softening within the resolution of the present measurements. The data for Sample 2 are somewhat noisy because of the small sample thickness. The softening in Sample 3 (type-Ib HPHT) is only approximately 1/3 of that of the Samples 1 and 2 (type IIa). This presumably is attributed to the neutralization of atomic vacancies by nitrogen doping. It is reasonable to assume that the concentration of neutral single vacancies in Sample 3 is less than in Samples 1 and 2 (type IIa). The inset of Fig. 3 shows data up to 100 K on a logarithmic temperature axis. Notably, the softening in Samples 2 and 3 is robust to external magnetic fields up to 14.0-16.5 T, similar as found for Sample 1 [20].
In summary, our ultrasound studies clearly reveal the presence of neutral single vacancies on the ppb level in all three diamond types investigated. In particular, we have uncovered that CVD diamonds, which are widely used for optical measurements in quantum technologies, also have atomic vacancies on the ppb level comparable to HPHT-grown diamonds. This result should be considered for future downsizing and removal of spin-noise sources in NV centers used as quantum devices. Importantly, the present study focused on atomic vacancies formed spontaneously during single-crystal growth, excluding those formed by neutron- or electron-beam irradiation. These results are of particular importance for researchers utilizing such diamonds in various fields, such as in quantum information, biological sensing, and power devices. Furthermore, our findings provide a novel perspective for studying atomic vacancies in diamonds. Further investigations of the elastic response of various types of diamonds at cryogenic temperature are required to explore in detail the underlying physics of diamond vacancies with respect to possible technological applications.
Acknowledgements.
We thank Profs. Terutaka Goto, Hiroaki Kusunose, and Satoru Hayami for helpful discussions. The present research was supported by JSPS KAKENHI Grants Nos. JP23H04868, JP21KK0046, JP22K03501, and Toyota Physical and Chemical Research Institute under the 2021 Toyota Riken Scholar Collaborative Research Program (Phase 1) to TY and TT. We acknowledge support from the DFG through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, project-id 390858490) and from the HLD at HZDR, a member of the European Magnetic Field Laboratory (EMFL). TY would like to acknowledge Kousuke Nakamura and Tatsuji Meike at Hokkaido University Technical Support Division for assistance in polishing the diamonds. TY and RH would like to thank Prof. Atsuhiko Miyata for supporting the measurements at HZDR.References
- Kasu et al. [2021] M. Kasu, N. C. Saha, T. Oishi, and S.-W. Kim, Fabrication of diamond modulation-doped FETs by NO delta doping in an AlO gate layer, Appl. Phys. Express 14, 051004 (2021).
- Zaitsev [2001] A. M. Zaitsev, Optical Properties of Diamond (Springer Berlin Heidelberg, 2001).
- Balmer et al. [2009] R. S. Balmer, J. R. Brandon, S. L. Clewes, H. K. Dhillon, J. M. Dodson, I. Friel, P. N. Inglis, T. D. Madgwick, M. L. Markham, T. P. Mollart, N. Perkins, G. A. Scarsbrook, D. J. Twitchen, A. J. Whitehead, J. J. Wilman, and S. M. Woollard, Chemical vapour deposition synthetic diamond: materials, technology and applications, J. Phys.: Cond. Matter 21, 364221 (2009).
- Sumiya et al. [2015] H. Sumiya, K. Harano, and K. Tamasaku, HPHT synthesis and crystalline quality of large high-quality (001) and (111) diamond crystals, Diamond and Related Materials 58, 221 (2015).
- Eeaton-Magaña and Shigley [2016] S. Eeaton-Magaña and J. E. Shigley, Observations on CVD-grown synthetic diamonds: A Review, Gems and Gemology 52, 222 (2016).
- Acosta and Hemmer [2013] V. Acosta and P. Hemmer, Nitrogen-vacancy centers: Physics and applications, MRS Bulletin 38, 127 (2013).
- Glenn et al. [2015] D. R. Glenn, K. Lee, H. Park, R. Weissleder, A. Yacoby, M. D. Lukin, H. Lee, R. L. Walsworth, and C. B. Connolly, Single-cell magnetic imaging using a quantum diamond microscope, Nat. Meth. 12, 736 (2015).
- Morishita et al. [2019] H. Morishita, T. Tashima, D. Mima, H. Kato, T. Makino, S. Yamasaki, M. Fujiwara, and N. Mizuochi, Extension of the Coherence Time by Generating MW Dressed States in a Single NV Centre in Diamond, Sci. Rep. 9, 13318 (2019).
- Barry [2020] J. F. Barry, Sensitivity optimization for NV-diamond magnetometry, Rev. Mod. Phys. 92, 015004 (2020).
- Hensen et al. [2015] B. Hensen, H. Bernien, A. E. Dréau, A. Reiserer, N. Kalb, M. S. Blok, J. Ruitenberg, R. F. L. Vermeulen, R. N. Schouten, C. Abellán, W. Amaya, V. Pruneri, M. W. Mitchell, M. Markham, D. J. Twitchen, D. Elkouss, S. Wehner, T. H. Taminiau, and R. Hanson, Loophole-free Bell inequality violation using electron spins separated by 1.3 kilometres, Nature 526, 682 (2015).
- Tashima [2019] T. Tashima, Experimental demonstration of two-photon magnetic resonances in a single-spin system of a solid, Phys. Rev. A 100, 023801 (2019).
- Cujia et al. [2019] K. S. Cujia, J. M. Boss, K. Herb, J. Zopes, and C. L. Degen, Tracking the precession of single nuclear spins by weak measurements, Nature 571, 230 (2019).
- Childress and Hanson [2013] L. Childress and R. Hanson, Diamond NV centers for quantum computing and quantum networks, MRS Bulletin 38, 134 (2013).
- Nakazato et al. [2022] T. Nakazato, R. Reyes, N. Imaike, K. Matsuda, K. Tsurumoto, Y. Sekiguchi, and H. Kosaka, Quantum error correction of spin quantum memories in diamond under a zero magnetic field, Commun. Phys. 5, 102 (2022).
- Rozpedek [2019] F. Rozpedek, Near-term quantum-repeater experiments with nitrogen-vacancy centers: Overcoming the limitations of direct transmission, Phys. Rev. A 99, 052330 (2019).
- Pezzagna et al. [2011] S. Pezzagna, D. Rogalla, D. Wildanger, J. Meijer, and A. Zaitsev, Creation and nature of optical centres in diamond for single-photon emission—overview and critical remarks, New J. Phys. 13, 035024 (2011).
- Lühmann et al. [2018] T. Lühmann, N. Raatz, R. John, M. Lesik, J. Rödiger, M. Portail, D. Wildanger, F. Kleißler, K. Nordlund, A. Zaitsev, J.-F. Roch, A. Tallaire, J. Meijer, and S. Pezzagna, Screening and engineering of colour centres in diamond, J. Phys. D: Appl. Phys. 51, 483002 (2018).
- Sumikura et al. [2020] H. Sumikura, K. Hirama, K. Nishiguchi, A. Shinya, and M. Notomi, Highly nitrogen-vacancy doped diamond nanostructures fabricated by ion implantation and optimum annealing, APL Materials 8, 031113 (2020).
- Luo et al. [2022] T. Luo, L. Lindner, J. Langer, V. Cimalla, X. Vidal, F. Hahl, C. Schreyvogel, S. Onoda, S. Ishii, T. Ohshima, D. Wang, D. A. Simpson, B. C. Johnson, M. Capelli, R. Blinder, and J. Jeske, Creation of nitrogen-vacancy centers in chemical vapor deposition diamond for sensing applications, New J. Phys. 24, 033030 (2022).
- [20] See Supplemental Material for (1) Thermodynamic Consideration of the Formation of Atomic Vacancies, (2) Neutral Single Vacancy V in Silicon and Diamond, (3) Experimental Details, (4) Ultrasonic Method, (5) Wave Functions for Hybrid Orbitals and Active Electric Quadrupole, (6) Quadrupolar Susceptibility, (7) Comparison of Si and diamond, (8) Magnetic Field Dependence of the Softening of , (9) Inconsistency of Quantum Ground State between Non-Irradiated and irradiated Diamond, and (10) Effect of NV, NV, P Centers, and Nuclear Spin.
- Bogdanov et al. [2018] K. V. Bogdanov, M. V. Zhukovskaya, V. Y. Osipov, E. V. Ushakova, M. A. Baranov, K. Takai, A. Rampersaud, and A. V. Baranov, Highly intensive emission of the NV centers in synthetic HPHT microdiamonds at low nitrogen doping, APL Materials 6, 086104 (2018).
- Stoneham [1977] A. Stoneham, The low-lying levels of the GR 1 centre in diamond, Solid State Commun. 21, 339 (1977).
- Subedi et al. [2021] S. Subedi, V. Fedorov, S. Mirov, and M. Markham, Spectroscopy of GR1 centers in synthetic diamonds, Optical Materials Express 11, 757 (2021).
- Ref [2022] Handbook of Scientific Tables (World Scientific, 2022).
- Coulson and Kearsley [1957] C. A. Coulson and M. J. Kearsley, Colour centres in irradiated diamonds. I, Proc. Royal Soc. Lond. A 241, 433 (1957).
- Yamada-Kaneta et al. [2006] H. Yamada-Kaneta, T. Goto, Y. Saito, Y. Nemoto, K. Sato, K. Kakimoto, and S. Nakamura, Vacancies in defect-free zone of point-defect-controlled CZ silicon observed by low-temperature ultrasonic measurements, Materials Science and Engineering: B 134, 240 (2006).
- Goto et al. [2006] T. Goto, H. Yamada-Kaneta, Y. Saito, Y. Nemoto, K. Sato, K. Kakimoto, and S. Nakamura, Observation of Low-Temperature Elastic Softening due to Vacancy in Crystalline Silicon, J. Phys. Soc. Jpn. 75, 044602 (2006).
- Goto et al. [2015] T. Goto, K. Mitsumoto, M. Akatsu, S. Baba, K. Okabe, R. Takasu, Y. Nemoto, H. Yamada-Kaneta, Y. Furumura, H. Saito, K. Kashima, and Y. Saito, Surface acoustic wave diagnosis of vacancy orbital with electric quadrupoles in silicon, J. Phys.: Conf. Ser. 592, 012150 (2015).
- Okabe et al. [2013] K. Okabe, M. Akatsu, S. Baba, K. Mitsumoto, Y. Nemoto, H. Yamada-Kaneta, T. Goto, H. Saito, K. Kashima, and Y. Saito, Strong Quadrupole-Strain Interaction of Vacancy Orbital in Boron-Doped Czochralski Silicon, J. Phys. Soc. Jpn. 82, 124604 (2013).
- Mitsumoto et al. [2014] K. Mitsumoto, M. Akatsu, S. Baba, R. Takasu, Y. Nemoto, T. Goto, H. Yamada-Kaneta, Y. Furumura, H. Saito, K. Kashima, and Y. Saito, Elastic Softening of Surface Acoustic Wave Caused by Vacancy Orbital in Silicon Wafer, J. Phys. Soc. Jpn. 83, 034702 (2014).
- Baba et al. [2011] S. Baba, T. Goto, Y. Nagai, M. Akatsu, H. Watanabe, K. Mitsumoto, T. Ogawa, Y. Nemoto, and H. Yamada-Kaneta, Quadrupole Effects of Vacancy Orbital in Boron-Doped Silicon, J. Phys. Soc. Jpn. 80, 094601 (2011).
- Goto et al. [2007] T. Goto, H. Yamada-Kaneta, K. Sato, M. Hikin, Y. Nemoto, and S. Nakamura, Observation of vacancy in crystalline silicon using low-temperature ultrasonic measurements, Physica B: Condensed Matter 401-402, 109 (2007).
- McSkimin and Andreatch [1972] H. J. McSkimin and P. Andreatch, Elastic Moduli of Diamond as a Function of Pressure and Temperature, Journal of Applied Physics 43, 2944 (1972).
- Migliori et al. [2008] A. Migliori, H. Ledbetter, R. G. Leisure, C. Pantea, and J. B. Betts, Diamond’s elastic stiffnesses from 322 K to 10 K, Journal of Applied Physics 104, 053512 (2008).
- Nagakubo et al. [2016] A. Nagakubo, M. Arita, H. Ogi, H. Sumiya, N. Nakamura, and M. Hirao, Elastic constant of C diamond between 10 and 613 K, Appl. Phys. Lett. 108, 221902 (2016).
- Laszlo et al. [2010] I. Laszlo, M. Kertesz, B. Slepetz, and Y. Gogotsi, Simulations of large multi-atom vacancies in diamond, Diamond and Related Materials 19, 1153 (2010).
- Kamihara et al. [2022] D. Kamihara, T. Shimizu, and K. Uchida, Discovery of Peculiar Electronic Structures of Decavacancy V in Silicon Crystal, J. Phys. Soc. Jpn. 91, 064709 (2022).
- Cox et al. [1994] A. Cox, M. E. Newton, and J. M. Baker, C,N and N ENDOR measurements on the single substitutional nitrogen centre (P1) in diamond, Journal of Physics: Condensed Matter 6, 551 (1994).
- Lüthi [2006] B. Lüthi, Physical Acoustics in the Solid State (Springer, Berlin, 2006).
- Clark and Walker [1973] C. Clark and J. Walker, The neutral vacancy in diamond, Proc. Royal Soc. Lond. A 334, 241 (1973).
- Davies and Penchina [1974] G. Davies and C. M. Penchina, The effect of uniaxial stress on the gr1 doublet in diamond, Proc. Royal Soc. Lond. A 338, 359 (1974).
- Nesheva [2023] D. Nesheva, Electron and neutron beam irradiation effects in homogeneous and nanostructured oxides, ACS Omega 8, 12603 (2023).
- Davies [1979] G. Davies, Dynamic jahn-teller distortions at trigonal optical centres in diamond, Journal of Physics C: Solid State Physics 12, 2551 (1979).
- Prentice [2017] J. C. A. Prentice, First-principles study of the dynamic jahn-teller distortion of the neutral vacancy in diamond, Physical Review B 95, 014108 (2017).
- Braukmann [2018] D. Braukmann, Circularly polarized zero-phonon transitions of vacancies in diamond at high magnetic fields, Physical Review B 97, 195448 (2018).
- Hayami [2018] S. Hayami, Classification of atomic-scale multipoles under crystallographic point groups and application to linear response tensors, Phys. Rev. B 98, 165110 (2018).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. Hollenberg, The nitrogen-vacancy colour centre in diamond, Physics Reports 528, 1 (2013).
- Lenef [1996] A. Lenef, Electronic structure of the N-V center in diamond: Theory, Phys. Rev. B 53, 13441 (1996).
- Schlüter [1985] M. Schlüter, Proc. Int. School of Physics “Enrico Fermi”, Proc. Int. School of Physics , p. 495 (1985).
- Yamada et al. [2010] T. Yamada, Y. Yamakawa, and Y. Ōno, Effect of the spin-orbit interaction and the electron-phonon coupling on the electronic state in a silicon vacancy, J. Phys.: Conf. Ser. 200, 012228 (2010).